Circular No. 8165

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2003 (283) 1 W. J. Merline, Southwest Research Institute (SwRI); C. Dumas, Jet

Propulsion Laboratory; N. Siegler and L. M. Close, University of Arizona; C. R. Chapman, P. M. Tamblyn, and D. Terrell, SwRI; A. Conrad, Keck Observatory; and F. Menard and G. Duvert, Observatoire de Grenoble, report the discovery on July 14.6 UT of a satellite of minor planet (283) Emma from K′-band imaging with the 10-m Keck II Telescope (+ NIRC2/AO adaptive optics system) on Mauna Kea. On July 14.5579 the satellite was at separation 0′′.26 (projected separation 370 km) and position angle 218^{o}; a detection in H-band was also made on July 14.6026. The satellite was also observed by the same team with the European Southern Observatory 8-m Very Large Telescope UT4/YEPUN (+ NAOS/CONICA adaptive optics system) on July 15.2750, 15.4229, and 16.4375 in bands J, H, and K_{s}. The H-band brightness difference is ∼ 5.5 mag, giving an estimated diameter of the satellite of ∼ 12 km.

NOVAE IN M31 Three additional apparent novae in M31 have been discovered, the first

two by K. Hornoch (cf. IAUC 8157), and the last two (independently) by M. Fiaschi, F. Di Mille, and D. Tiveron (cf. IAUC 8155). The tabulated data for the first object are from Hornoch; the remaining tabulated data are from Fiaschi et al. (whose magnitudes are with an Hα filter). The second object was found independently by Hornoch near the limit of seven CCD images, and he provides position end figures 41^{s}.0, 33′′ (with offset 43′′ west and 142′′ north of the center of M31).

2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
July 14.99 0^{h}42^{m}15.^{s}81 +41^{o}12′00′′.5 16.9 318′′ W, 252′′ S
July 19.986 0 42 41.10 +41 18 32.0 15.9 36′′ W, 144′′ N
July 19.986 0 42 45.10 +41 17 53.7 14.9 9′′ E, 105′′ N

Additional [Hα] magnitudes by Fiaschi et al.: Second object, July 15 UT, [18; 21.020, 16.0. Third object, July 15, [18; 21.020, 15.0. Further approximate R-band magnitudes by Hornoch: First object, July 11.04, [18.7; 15.99, 17.5; 19.006, 17.5. Second object, July 14.988, [18.8; 19.031, 18.5; 21.00, 18.9. Additional R magnitudes by Hornoch, from images taken by himself and by L. Sarounova, of the apparent nova reported on IAUC 8157: June 30.022, 18.3; July 2.997, 18.6; 3.034, 18.7; 7.993, 18.0; 9.011, 18.7.

2003 July 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green